OpenStudy (anonymous):

How to solve trig equation ? \( tan^2=4 \)

OpenStudy (anonymous):

Do I just \( \sqrt{tan^2} = \sqrt{4}\) \( \sqrt{tan^2} = \sqrt{4}\) \( tan = 2\)

OpenStudy (amoodarya):

you don't forget something ? \[tanx=\pm2\]

OpenStudy (anonymous):

Ok, so I was on the right path